Hoe much soil containing 45% sand do you need to add to 1ft(cubed) of soil containing 15% sand in order to make a soil containing 35% sand?
Answer by stanbon(75887)

How much soil containing 45% sand do you need to add to 1ft(cubed) of soil containing 15% sand in order to make a soil containing 35% sand?
-------------------
Equation:
sand + sand = sand
0.45x + 0.15*1 = 0.35(x+1)
----
45x 15*1 = 35x + 35*1
-------
10x = 20
x = 2 cu. ft (amt of 45% soil needed)
=======================================
Cheers,
Stan H.